# Thermodynamics of black and white holes in ensemble of Planckons

### v3

@@ II.2 Entropy in terms of Planck-scale elements
In Eq.(6), the entropy of the ensemble arises from the correlations between Plancons. It can be interpreted as the number of $\cal N$ ways to select two Plancons from an ensemble of $N$ Plancons:
@@ II.3 Splitting of black hole into smaller parts and back reaction
$w(N\rightarrow 1+1+...+1)=e^{-\cal N}=\exp\left(-\frac{N(N-1)}{2}\right)=\exp(-S_{\rm BH}(N))\,.$ (15)
@@
$w(N\rightarrow 1+1+...+1)=\Pi_{k=1}^{N}e^{-(N-k)}=\exp\left(-\frac{N(N-1)}{2}\right)=\exp(-S_{\rm BH}(N))\,.$ (16)
@@ III.1 Probability of formation of white hole and negative entropy
$w(1+1+...+1\rightarrow N)=\exp\left(-\frac{N(N-1)}{2}\right)\,,$ (17)
This means that formation of the white hole by merging of Planckons represent the rare thermodynamic fluctuation, at which the entropy decreases from $S=0$ to the maximum negative value, which corresponds to the entropy of a black hole with a minus sign:
@@
The entropy of a white hole is the negative entropy of the least probable exceptional state. The white hole relaxes with increasing entropy, and this corresponds to the increase in entropy during the formation of a black hole.
@@
In the same manner the time reversal transforms the expanding de Sitter with the positive Hubble parameter, $H>0$, to contracting de Sitter with $H<0$. Correspondingly, due to anti-symmetry of the Hubble parameter, $TH=-H$, the entropy of the volume inside the cosmological horizon is negative in the contracting Universe, $S(H<0)=-S(H>0)=-A/4G$, where $A$ is the area of the cosmological horizon.Volovik2025b
@@ IV.3 Horizons, thermodynamics and Lifshitz transition
The equation (28) is also applicable to the processes associated with the cosmological horizon. The corresponding temperature governing the processes of radiation from the horizon is the Gibbons-Hawking temperature $T_{\rm GH}=H/2\pi$, where $H$ is the Hubble parameter. The local temperature of the de Sitter state, describing the de Sitter thermodynamics, is not related to horizon and is twice as large.Volovik2025b Nevertheless, the entropy of the Hubble volume (the volume inside the cosmological horizon) corresponds to the Gibbons-Hawking entropy of the cosmological horizon, $S_{H}=A/4G$, where $A$ is the area of the cosmological horizon. This connection between local thermodynamics and the thermodynamics of the cosmological horizon is a consequence of the specific symmetry of the de Sitter state and is supported by topological arguments.Kusmartsev2025 ; Volovik2025f

### v4

@@ II.2 Entropy in terms of Planck-scale elements
In Eq.(6), the entropy of the ensemble arises from the correlations between Plancons. It can be interpreted as the number $\cal N$ of ways to select two Planckons from an ensemble of $N$ Planckons:
@@ II.3 Splitting of black hole into smaller parts and back reaction
$w(N\rightarrow\underbrace{1+1+...+1}_{N}\,)=e^{-\cal N}=\exp\left(-\frac{N(N-1)}{2}\right)=\exp(-S_{\rm BH}(N))\,.$ (15)
@@
$w(N\rightarrow\underbrace{1+1+...+1}_{N}\,)=\Pi_{k=1}^{N}e^{-(N-k)}=\exp\left(-\frac{N(N-1)}{2}\right)=\exp(-S_{\rm BH}(N))\,.$ (16)
@@ III.1 Probability of formation of white hole and negative entropy
$w(\,\underbrace{1+1+...+1}_{N}\,\rightarrow N)=\exp\left(-\frac{N(N-1)}{2}\right)=e^{-\cal N}\,.$ (17)
This means that formation of the white hole by merging of Planckons represents the rare thermodynamic fluctuation, at which the entropy decreases from $S=0$ to the maximum negative value, which corresponds to the entropy of a black hole with a minus sign:
@@
This equation suggests the analog of Bekenstein entropy boundBekenstein1981 ; Bekenstein2004 and holographic entropy boundHooft1993 ; Susskind1995 , which now includes the negative entropy of white holes. The entropy of an ensemble of $N$ Planckons is bounded by maximum and minimum values that correspond to the entropies of black and white holes:
$-\frac{N(N-1)}{2}\leq S(N)\leq\frac{N(N-1)}{2}\,.$ (19)
The entropy of a white hole is the negative entropy of the least probable exceptional state. A white hole relaxes into a black hole with increasing entropy, and thus due to the negative entropy of a white hole, the second law of thermodynamics is not violated.
@@
In the same manner the time reversal transforms the expanding de Sitter with the positive Hubble parameter, $H>0$, to contracting de Sitter with $H<0$. Correspondingly, due to anti-symmetry of the Hubble parameter, $TH=-H$, the entropy of the volume inside the cosmological horizon (the Hubble volume) is negative in the contracting Universe, $S(H<0)=-S(H>0)=-A/4G$, where $A$ is the area of the cosmological horizon.Volovik2025b In transition between expanding and contracting Universe via Minkowskii state, the local entropy density $s$ changes sign by crossing zero value, which is natural for the empty Minkowski vacuum. But the total entropy $S(H)=sV_{H}$ of the Hubble volume $V_{H}$ changes sign by crossing an infinite value. This shows that spatial infinity in the Minkowski vacuum can play a role similar to that of the cosmological event horizon. Escape to infinity may play the same role as escape through the crossing of the cosmological horizon.
@@ IV.3 Horizons, thermodynamics and Lifshitz transition
The equation (29) is also applicable to the processes associated with the cosmological horizon. The corresponding temperature governing the processes of radiation from the horizon is the Gibbons-Hawking temperature $T_{\rm GH}=H/2\pi$, where $H$ is the Hubble parameter. The local temperature of the de Sitter state, describing the de Sitter thermodynamics, is not related to horizon and is twice as large.Volovik2025b Nevertheless, the local entropy density $s$ integrated over the Hubble volume $V_{H}$ (the volume inside the cosmological horizon) corresponds to the Gibbons-Hawking entropy of the cosmological horizon, $S_{H}=sV_{H}=A/4G$, where $A$ is the area of the cosmological horizon. This connection between local thermodynamics and the thermodynamics of the cosmological horizon is a consequence of the specific symmetry of the de Sitter state and is supported by topological arguments.Kusmartsev2025 ; Volovik2025f